# RoboDuet: A Framework Affording Mobile-Manipulation and Cross-Embodiment

### v1

@@ III Methods
Figure 1: RoboDuet consists of two stages. In Stage 1, the legged robot is trained to learn a loco policy with arm fixed. The loco policy inputs body orientation commands, proprioceptive states, and privileged information, and outputs action $a_{t}^{loco}$ for each joint to achieve specific loco goals. In Stage 2, the loco policy and the arm policy are trained simultaneously. The loco policy from Stage 1 is used, taking the same inputs while the body orientation commands are provided by the arm policy. The arm policy is trained to achieve given 6D end-effector poses by taking as inputs the target pose and other observations and outputting actions for each joint, $a_{t}^{arm^{J}}$, as well as body orientation commands for the locomotion policy.
@@ III-B Two stage training
Stage 2 aims to coordinate locomotion and manipulation to achieve whole-body large-range mobile manipulation. The arm policy will be activated simultaneously with all the robotic arm joints. We adopt 6D spatial target pose of end-effector as policy input. To eliminate the influence brought by body rotation [8], we similarly use a posture-independent spherical coordinate to define the target end-effector pose $\chi_{t}$. The target end-effector position is represented by radius $l$, latitude $p$, and longitude $y$. To improve the accuracy of end-effector orientation tracking, we use Euler angles in Z-X-Y order for sampling, which can intuitively exclude many illegal postures, and convert them to included angle along each axis of the coordinate. The mathematical form of sampling can be expressed as follows:
@@ IV-B1 Experiment Setup
To validate the significance of the two-stage training and the cooperative policy, which are key components of RoboDuet, we establish a Baseline algorithm combining one-stage training with a unified policy. The Two-Stage algorithm modifies this baseline by transitioning from one-stage to two-stage training, while the Cooperated algorithm builds on the baseline by replacing the unified policy with a cooperative policy. RoboDuet itself incorporates both two-stage training and cooperative policy. We train all algorithms for 45,000 iterations across 3 seeds, with the two-stage training comprising 10,000 iterations for stage 1 and 35,000 for stage 2, keeping the rest training components constant. This process generates the corresponding policies for each algorithm.
@@ IV-C1 Experiment Setup
Furthermore, we expand the application of RoboDuet by incorporating a different robotic arm, the WidowX 250s, with each of the three legged robots. This process demonstrates the framework’s versatility, successfully training each combination. The effectiveness of these configurations is illustrated in the bottom row of Fig. LABEL:fig:setup, where they are depicted in various poses, showcasing the robust training outcomes across diverse robot embodiments.
@@ IV-D Real-World Deployment
In terms of real-world deployment, we craft three distinct types of tasks to validate the effectiveness of our policy in the real world and its proficiency in handling diverse mobile manipulation challenges. We deploy the RoboDuet policy in the real world on Go1+Arx5 and control the robot system in two different ways. The first invloves utilizing Apriltag to obtain the distance between the target end-effector pose and the current pose, and the other employs the controller to pre-define the target pose relative to the base coordinate of the legged robot.
@@ V Discussion and Limitations
Due to the small field of view (FOV) of the RealSense D435i, it is challenging to get a complete view of the AprilTag when the end-effector approaches the target position. This makes using the AprilTag to obtain the target pose less robust. We have tried using AnyGrasp to infer the object pose, but its inference frequency is far below the desired 50Hz, and it cannot run onboard, making it unsuitable for the trained policy. Since this work mainly focuses on RL-based whole-body control, we have not further attempted to integrate vision-related content. In the future, we can try connecting state-of-the-art pose estimation algorithm and navigation module to achieve truly universal mobile manipulation.

### v2

@@ III Methods
Figure 1: RoboDuet consists of two stages. In Stage 1, the legged robot is trained to learn a loco policy with arm fixed. The loco policy inputs body orientation commands, proprioceptive states, and privileged information, and outputs action $a_{t}^{loco}$ for each joint to achieve specific loco goals. In Stage 2, the loco policy and the arm policy are trained simultaneously. The loco policy from Stage 1 is used, taking the same inputs while the body orientation commands are provided by the arm policy. The arm policy is trained to achieve given 6D poses by taking the target end-effector pose and other observations as input and outputting actions $a_{t}^{arm^{J}}$ for each joint, as well as body orientation commands for the locomotion policy.
@@ III-B Two stage training
Stage 2 aims to coordinate locomotion and manipulation to achieve whole-body large-range mobile manipulation. The arm policy will be activated simultaneously with all the robotic arm joints. We adopt 6D spatial target pose of end-effector as policy input. To eliminate the influence brought by body rotation [8], we similarly use a posture-independent spherical coordinate to define the target end-effector pose $\chi_{t}$. The target position of end-effector is represented by radius $l$, latitude $p$, and longitude $y$. To improve the accuracy of end-effector orientation tracking, we use Euler angles in Z-X-Y order for sampling, which can intuitively exclude many illegal postures, and convert them to included angle along each axis of the coordinate. The mathematical form of sampling can be expressed as follows:
@@ IV-B1 Experiment Setup
To validate the significance of the two-stage training and the cooperative policy, which are key components of RoboDuet, we establish a Baseline algorithm training a unified policy in one-stage. The Two-Stage algorithm modifies this baseline by transitioning from one-stage to two-stage training, while the Cooperated algorithm builds on the baseline by replacing the unified policy with a cooperative policy. RoboDuet itself incorporates both two-stage training and cooperative policy. We train all algorithms for 45,000 iterations across 3 seeds, with the two-stage training comprising 10,000 iterations for stage 1 and 35,000 for stage 2, keeping the rest training components constant. This process generates the corresponding policies for each algorithm.
@@ IV-C1 Experiment Setup
Furthermore, we expand the application of RoboDuet by incorporating a different robotic arm, the WidowX 250s, with each of the three legged robots. This process demonstrates the framework’s versatility, successfully training each combination. Furthermore, we expand the application of RoboDuet by incorporating WidowX 250s with each of the three legged robots, underscoring the system’s flexibility. This integration allows us to demonstrate that each combination can be efficiently trained using the same set of hyperparameters, minimizing the need for extensive adjustments. The effectiveness of these configurations is illustrated in the top row of Fig. LABEL:fig:setup, where they are depicted in various poses, showcasing the robust training outcomes across diverse robot embodiments.
@@ IV-D Real-World Deployment
In terms of real-world deployment, we craft three distinct types of tasks to validate the effectiveness of our policy in the real world and its proficiency in handling diverse mobile manipulation challenges. We deploy the RoboDuet policy in the real world on Go1+Arx5 and control the robot system in two different ways. The first invloves utilizing camera to obtain the distance between the target end-effector pose and the current pose in real time, and the other employs the controller to pre-define the target end-effector pose relative to the base coordinate of the legged robot.
@@ V Discussion and Limitations
Due to the small field of view (FOV) of the RealSense D435i, it is challenging to get a complete view of the AprilTag when the end-effector approaches the target position. This makes using the AprilTag to obtain the target end-effector pose less robust. We have tried using AnyGrasp to infer the object pose, but its inference frequency is far below the desired 50Hz, and it cannot run onboard, making it unsuitable for the trained policy. Since this work mainly focuses on RL-based whole-body control, we have not further attempted to integrate vision-related content. In the future, we can try connecting state-of-the-art pose estimation algorithm and navigation module to achieve truly universal mobile manipulation.